A homeowner can obtain a $250,000, 30-year fixed-rate mortgage at a rate of 6.0 percent with zero points or at a rate of 5.5 percent with 2.25 points. How long must the owner stay in the house to make it worthwhile to pay the points if the payment saving is invested monthly?

Question 8 options:

A)

6.04 years

B)

7.15 years

C)

3.33 years

D)

5.90 years

E)

more than 30 years
Answer #1

Answer is

D)

5.90 years


N=360
Loan A = $250,000

I/y = 6/12

FV=0

PMT = $1498.88

Loan B = $250,000

N=360

I/y= 5.5/12

FV=0

PMT = $1,419.47

Savings = $79.40(Loan A – Loan B)

Points = $250000*2.25% =$5625

Break even point = points / savings = 5625/79.40 =70.84 months = 70.84/12 =5.90 years